OpenStudy (anonymous):

In what type of regular polygon is each interior angle equal to 168°?

OpenStudy (amistre64):

an interior angle is the sum of two adjacent base angles of an iso triangle. 180 - 168 = the measure of its central angle

OpenStudy (amistre64):

the central angle is a portion of the whole circle; and when we divide 360 by the central angle it tells us how many sides we got

OpenStudy (amistre64):

if me head is on straight, i get a central angle of 12; and 360/12 = the number of sides

OpenStudy (anonymous):

Exterior is 12, 360/12

OpenStudy (anonymous):

Exterior angles sum to 360.

OpenStudy (amistre64):

is exterior the sweep of the outside corner? or the sweep made by the supplement to the interior?

OpenStudy (anonymous):

12 +168 =180

OpenStudy (anonymous):

Actually it is true for any convex polygon, not just regular ones.

OpenStudy (amistre64):

http://mathworld.wolfram.com/ExteriorAngle.html

OpenStudy (anonymous):

My bad terminology:-( Answer is right though...

OpenStudy (amistre64):

yes, answer is right :)

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Polygon says exterior is the way I had it. We all know Wikipedia is unreliable, though:-)

OpenStudy (anonymous):

Actually, searching around, it is Wolfram's definition that is at variance with most other sources, so I think I will stick to my terminology after all:-)